OpenStudy (anonymous):

f'((x^2 + 1)/x) im not sure what to do about the denominator. are you supposed to transform it so its not a fraction?

OpenStudy (kamibug):

could you please use the equation tool to better understand that? =)

OpenStudy (anonymous):

\[\prime \frac{ x ^{2} +1}{ x}\] its supposed to f prime but it doesnt show up well

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{x^2+1}{x}\right)'\]I'm not sure why there is an f there... Are you just trying to take the derivative of this?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

We'll need to apply the quotient rule:\[\Large\bf\sf \left(\frac{u}{v}\right)'\quad=\quad \frac{\color{royalblue}{\left[u\right]'}v-u\color{royalblue}{\left[v\right]'}}{v^2}\]Taking the derivative of the blue parts.\[\Large\bf\sf \left(\frac{x^2+1}{x}\right)'\quad=\quad \frac{\color{royalblue}{\left[x^2+1\right]'}x-(x^2+1)\color{royalblue}{\left[x\right]'}}{x^2}\]

zepdrix (zepdrix):

Understand the setup?

OpenStudy (anonymous):

yeah, its just that the section the problem is in is a few before they teach the quotient rule. i was trying to see if there was another way to do it algebraically or if you could separate transform it into exponents like \[\left( x ^{-1} \right)\left( x ^{2} +1 \right)\]